OpenStudy (londoncat):

the roots of the quadratic equation z^2 + az + b = 0 are 2 - 3i and 2 + 3i. What is a+b?

OpenStudy (campbell_st):

ok... so if you use the roots, you should know \[x = 2 \pm \sqrt{-9}\] so this becomes \[x - 2 = \sqrt{-9}\] so you should be able to get the original quadratic from here

OpenStudy (londoncat):

but how do you get a b? im sorry im confused

OpenStudy (londoncat):

@campbell_st

OpenStudy (campbell_st):

ok... let's just work on getting the factored quadratic so is you square both sides you get \[(z^2 -2)^2 = -9\] or \[z^2 - 4z + 4 = -9\]

OpenStudy (campbell_st):

so add 9 to both sides and compare it to the original to find a and b then you can find a + b

OpenStudy (londoncat):

z^2-4x+13=0?

OpenStudy (campbell_st):

great so a=? and b = ? then find a + b

OpenStudy (londoncat):

4 and 13

OpenStudy (campbell_st):

almost.. check the sign for a...

OpenStudy (londoncat):

so 17 is a+b?

OpenStudy (campbell_st):

well just check \[z^2 + az + b = 0\] \[z^2 - 4z + 13 = 0\]

OpenStudy (londoncat):

wait so is it -4

OpenStudy (campbell_st):

correct

OpenStudy (londoncat):

so total is 9

OpenStudy (campbell_st):

correct

OpenStudy (londoncat):

thanks

OpenStudy (sshayer):

\[ax^2+bx+c=0\] \[\sum~of~roots=-\frac{ b }{ a }\] product of roots=c/a